OpenStudy (anonymous):

Solve for x. -x + 4 = 32 28?

OpenStudy (anonymous):

it has a negative sign on it. It's -28.

OpenStudy (anonymous):

now move the negative sign

OpenStudy (shamil98):

-x + 4 = 32 -x = 28 you need to divide by -1 to get x by itself x =- 28